Find the mean and standard error of the sample means that is normally distributed with a mean of 10 and a standard deviation of 2 and a sample size of 25.
Using the above distribution find the probability of a score less than 11.

Answer by stanbon(75887)

Find the mean and standard error of the sample means that is normally distributed with a mean of 10 and a standard deviation of 2 and a sample size of 25.
-----
Ans:: mean of the sample means = mean of the population = 10
std of the sample means = (std of population)/sqrt(n) = 2/sqrt(25) = 2/5
----------------------------------------------------
Using the above distribution find the probability of a score less than 11
Assuming that 11 is a score and not a sample mean::
z(11) = (11-10)/2 = 1/2
P(x < 11) = P(z < 1/2) = 0.6915
-----
Assuming that 11 is sample mean::
z(11) = (11-10)/[2/sqrt(25)] = 1/[2/5] = 2.5
P(x < 11) = P(z < 2.5) = 0.9938
==================
Cheers,
Stan H.
------------------